OpenStudy (anonymous):

The function W=Z^4-Z^3/3 has a(n); 1.minimum at Z=0 and a maximum at Z=¼ 2.minimum at Z=0 and a point of inflection at Z=¼ 3.maximum at Z=0 and a minimum at Z=¼ 4.point of inflection at Z=0 and a minimum at Z=1/4

OpenStudy (anonymous):

dw/dz= 4z^3-z2=0 z=0, z=1/4. w''= 12z^2-2z=0, when we plug z=0, and negative when plug z=1/4. maximum at z=1/4. and neither maximum nor minimum at z=0